# A study of bound states of $\eta,\eta^{{}^{\prime}},D^{0}$, $\bar{D^{0}}$, $B^{0}$, $\bar{B^{0}}$, $\bar{K^{0}}$ and $\phi$ mesons with light and heavy nuclei within chiral SU(3) model

### v1

@@ I Introduction
In the present paper, we shall explore the formation of bound states of pseudoscalar mesons, $\eta,\eta^{{}^{\prime}},D^{0},\bar{D^{0}},B^{0},\bar{B^{0}},\bar{K^{0}}$ and the vector meson $\phi$, whereas, the bound states of $\omega$, $J/\psi$, $\eta_{c}$ and $\eta_{b}$ will be investigated in future work. For our current work, we shall use the chiral SU(3) hadronic mean field model which is based on low energy properties of QCD [63]. Chiral SU(3) model has been used in literature to study the modification of properties of kaons [65, 66, 64], vector mesons ($\omega,\rho$ and $\phi$) [67], pseudoscalar $\eta$ meson [68, 69], $D$ [38, 70, 39] and $B$ mesons [71], charmonium [72] and bottomonium [73], in nuclear and strange hadronic matter. The optical potentials of mesons under study, calculated using chiral SU(3) for different light and heavy nuclei, will be used to solve Klein Gordon equation in the momentum space to obtain the binding energies and decay width.
@@ II Chiral SU(3) hadronic model
To investigate the formation of mesic nuclei in the present work, we shall use the chiral SU(3) hadronic mean-field model based on nonlinear realization of chiral symmetry [63, 74, 75, 76]. In the chiral SU(3) model, the interactions between the nucleons are mediated through the exchange of non-strange scalar meson $\sigma$, strange scalar meson $\zeta$, scalar-isovector meson $\delta$, vector meson $\omega$ and vector-isovector meson $\rho$. The scalar field $\delta$ and vector field $\rho$ contribute when medium has finite isospin asymmetry. In this model, the scalar dilaton field, $\chi$, known as glueball field is introduced to incorporate the broken scale invariance property of the QCD [63]. The Lagrangian density for chiral SU(3) hadronic mean-field model is written as [63]
@@
respectively. In the above, $\gamma_{i}$ is the degeneracy factor for the nucleons.
@@ III.1 Pseudoscalar $\eta$ and $\eta^{{}^{\prime}}$ mesons
$m_{\eta_{0}\eta_{8}}^{*}=\left[\frac{1}{\sqrt{3}\sigma_{0}c}\left(m_{\pi}^{2}f_{\pi}-m_{K}^{2}f_{K}\right)\right]^{1/2}.$ (38)
@@ III.2 Pseudoscalar $D^{0}$ and $\bar{D^{0}}$ mesons
The effective masses of neutral $D^{0}$ and $\bar{D^{0}}$ mesons required in Eq.(22) to study the $D$ mesic nuclei are obtained in the present work by extending the chiral SU(3) model to the SU(4) case [38]. The mesons $D^{0}$ and $\bar{D^{0}}$ are the members of $D$ $\left(D^{+},D^{0}\right)$ and $\bar{D}$ $\left(D^{-},\bar{D^{0}}\right)$ doublets, respectively. From the Lagrangian density describing the interactions of $D$ and $\bar{D}$ mesons with nucleons in the nuclear medium, equations of motion are obtained whose Fourier transformation will give the dispersion relation
@@ III.3 Pseudoscalar $B^{0}$ and $\bar{B^{0}}$ mesons
To study the formation of bound states corresponding to $B^{0}$ and $\bar{B^{0}}$ mesons chiral SU(3) model will be generalized to SU(5) sector as these bottom mesons are composed of one heavy $b$ quark/antiquark. The pseudoscalar $B^{0}$ and $\bar{B^{0}}$ mesons belong to the open bottom $B(B^{+},B^{0})$ and $\bar{B}(B^{-},\bar{B^{0}})$ meson doublets. In Ref. [43] the masses of open bottom mesons have been calculated in nuclear and strange hadronic matter using chiral effective model. For our present calculations, we need the mass shift of $B^{0}$ and $\bar{B^{0}}$ mesons in the nuclear medium at zero temperature obtained by solving the dispersion relation
@@
$+\left[-\frac{1}{f_{B}}\left(\sigma^{\prime}+\sqrt{2}\zeta_{b}^{\prime}+\delta^{\prime}\right)+\frac{d_{1}}{2f_{B}^{2}}\left(\rho_{p}^{s}+\rho_{n}^{s}\right)\right.\left.+\frac{d_{2}}{2f_{B}^{2}}\left(\rho_{p}^{s}+2\rho_{n}^{s}\right)\right]\left(\omega^{2}-\textbf{k}^{2}\right),$ (46)
@@
$+\left[-\frac{1}{f_{B}}\left(\sigma^{\prime}+\sqrt{2}\zeta_{b}^{\prime}+\delta^{\prime}\right)+\frac{d_{1}}{2f_{B}^{2}}\left(\rho_{p}^{s}+\rho_{n}^{s}\right)\right.\left.+\frac{d_{2}}{2f_{B}^{2}}\left(\rho_{p}^{s}+2\rho_{n}^{s}\right)\right]\left(\omega^{2}-\textbf{k}^{2}\right),$ (47)
@@ III.4 Pseudoscalar $\bar{K^{0}}$ mesons
$\Pi_{\bar{K^{0}}}(\omega,|\textbf{k}|,\rho(r))$ $=\frac{1}{2f_{K}^{2}}\left(\rho_{p}+2\rho_{n}\right)\omega+\frac{m_{B}^{2}}{2f_{B}}\left(\sigma^{\prime}+\sqrt{2}\zeta^{\prime}-\delta^{\prime}\right)$
$+\left[-\frac{1}{f_{K}}\left(\sigma^{\prime}+\sqrt{2}\zeta^{\prime}+\delta^{\prime}\right)+\frac{d_{1}}{2f_{K}^{2}}\left(\rho_{p}^{s}+\rho_{n}^{s}\right)\right.\left.+\frac{d_{2}}{2f_{K}^{2}}\rho_{n}^{s}\right]\left(\omega^{2}-\textbf{k}^{2}\right).$ (49)
@@ IV Results and discussion
We now discuss the possibility of the formation of mesic nuclei of various mesons through the numerical calculations of the optical potentials and binding energies. We shall calculate the optical potentials of $\eta,\eta^{{}^{\prime}}$, $D^{0}$, $\bar{D^{0}}$, $B^{0}$, $\bar{B^{0}}$,$\bar{K^{0}}$ and $\phi$ mesons in the medium relevant for the nuclei 12C, 16O, 40Ca and 208Pb. As we discussed in the previous section, the in-medium properties of mesons are calculated using the chiral SU(3) hadronic mean field model and extended to SU(4) and SU(5) sector to obtain the in-medium masses of charmed and bottom mesons [63].
@@
Figure 1: The mass shift of $\eta$ [in subplot (a)] and $\eta^{\prime}$ [in subplot (b)] mesons are shown as a function of density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of nuclear medium for isospin asymmetry, $I=0,0.211$ and $0.3$.
@@ IV.1 Binding energies for $\eta$ and $\eta^{{}^{\prime}}$ mesic nuclei
In this subsection, we shall explore the possibility of the formation of $\eta$ and $\eta^{{}^{\prime}}$ mesic nuclei. Using Eq. (39), the mass shifts of pseudoscalar $\eta$ and $\eta^{{}^{\prime}}$ mesons are calculated within the chiral SU(3) model. In Fig. 1, we have shown the mass shift of $\eta$ and $\eta^{{}^{\prime}}$ mesons as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of the infinite nuclear medium for isospin asymmetry $I=0,0.211$ and $0.3$. The isospin asymmetry $I=0.211$ corresponds to the heavy nuclei 208Pb considered in the present work for the bound state formation. As we discussed earlier, for the calculations of $\eta$ and $\eta^{{}^{\prime}}$ meson properties, we considered the mixing of $\eta_{8}$ and $\eta_{0}$ states. The pseudoscalar $\eta$ meson which is member of octet is observed to undergo larger mass shift as compared to the singlet $\eta^{{}^{\prime}}$. At nuclear saturation density $\rho_{B}=\rho_{0}$, for isospin asymmetry $I=0(0.3)$, the values of mass shift $\Delta m_{\eta}$ and $\Delta m_{\eta}^{{}^{\prime}}$ are observed to be $-34.64(-34.21)$ and $-1.03(-1.02)$ MeV, respectively. At baryon density $\rho_{B}=4\rho_{0}$, the above values of mass shift change to $-124.15(-120.92)$ and $-3.00(-2.94)$ MeV. The finite isopsin asymmetry of the medium causes less drop in the masses of $\eta$ and $\eta^{{}^{\prime}}$, although the impact is small. Furthermore, in the present chiral model calculations, compared to $\eta$ mesons, the singlet $\eta^{{}^{\prime}}$ undergoes very less drop in mass as a function of density of nuclear medium.
@@ IV.2 Binding energies for $D^{0}$ and $\bar{D^{0}}$ mesic nuclei
Now we discuss the formation of bound states for neutral $D^{0}$ and $\bar{D^{0}}$ mesons with the nuclei within the chiral SU(3) hadronic mean field model. The mass shift of $D^{0}$ and $\bar{D^{0}}$ mesons calculated using Eq. (44) are plotted in Fig. (3) as a function of baryon density ratio $\rho_{B}/\rho_{0}$. As can be seen from the figure, the in-medium masses of both $D^{0}$ and $\bar{D^{0}}$ mesons decreases as a function of density of the nuclear medium. Furthermore, the increase of isospin asymmetry of the medium from zero to finite value (which is relevant for the study of nuclei having different number of protons and neutrons) decreases the magnitude of mass shift for these neutral pseudoscalar mesons. At baryon density $\rho_{B}=\rho_{0}(4\rho_{0})$, the values of mass shift for $D^{0}$ and $\bar{D^{0}}$ mesons are observed to $-76.96(-343.82)$ and $-26(-159.71)$ MeV in the symmetric nuclear medium, whereas for $\eta=0.3$ these values changes to $-64.17(-298.80)$ and $-23.37(-148.43)$ MeV, respectively. Using the observed values of mass shift for $D^{0}$ and $\bar{D^{0}}$ mesons at center of given nuclei in Eq. (22), the real part of optical potential is calculated and is plotted in Fig. 4 as a function of $r$ for nuclei, 12C, 16O, 40Ca and 208Pb. As one can see from the figure, depth of optical potential (large negative value) is observed to be more for $D^{0}$ mesons as compared to $\bar{D^{0}}$. Furthermore, the imaginary part of optical potential, required to understand the absorption of mesons in the nuclei and its impact on the binding energy of neutral $D$ mesic nuclei, is calculated using Eq. (23). The values of binding energy and decay width obtained by solving the Klein Gordon equation are given in Tables 2 and 3, for $D^{0}$ and $\bar{D^{0}}$ mesons, respectively, for $\kappa=0,0.5$ and 1. As one can see from Table 2, the bound states of $D^{0}$ mesons, for 40Ca and 208Pb may be found not only for lower states $1s$ and $1p$, but also for other higher states when $\kappa=0$ and $0.5$. When the value of $\kappa$ is increased to one, the magnitude of binding energy is observed to be less as compared to the decay width value. This may cause a broadening of the peak and hindrance in the observation of clear signature of bound states. Comparing the values of binding energies in Tables 2 and 3, magnitude of binding energies for $\bar{D^{0}}$ mesons are significantly less than $D^{0}$ mesons. In Ref. [10], the bound states of $D^{0}$ mesons have been studied for several nuclei using the coupled channel approach. Comparing our results for 12C, 40Ca and 208Pb with Ref. [10], the magnitude of binding energies are observed to be more in our case. The binding energies of $D^{0}$ and $\bar{D^{0}}$ mesons for 208Pb nuclei are calculated using QMC model and are found to approximately $-96$ and $-25$ MeV, respectively [7].
@@
Figure 3: The mass shift of $D^{0}$ [in subplot (a)] and $\bar{D^{0}}$ [in subplot (b)] mesons are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of nuclear medium for isospin asymmetry, $I=0,0.211$ and $0.3$.
@@
Figure 4: The real optical potential $U(r)$ of $D^{0}$ [in subplot (a)] and $\bar{D^{0}}$ [in subplot (b)] mesons are shown as a function of distance $r$ from center, for nuclei 12C, 16O, 40Ca and 208Pb, with mass number, $A=12,16,40$ and $208$, respectively.
@@
Figure 5: The mass shift of $B^{0}$ [in subplot (a)] and $\bar{B^{0}}$ [in subplot (b)] mesons are shown as a function of density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of nuclear medium for isospin asymmetry, $I=0,0.211$ and $0.3$.

### v2

@@ I Introduction
In the present paper, we shall explore the formation of bound states of pseudoscalar mesons, $\eta,\eta^{{}^{\prime}},D^{0},\bar{D^{0}},B^{0},\bar{B^{0}},\bar{K^{0}}$ and the vector meson $\phi$, whereas, the bound states of $\omega$, $J/\psi$, $\eta_{c}$ and $\eta_{b}$ will be investigated in future work. For our current work, we shall use the chiral SU(3) hadronic mean field model which is based on low energy properties of QCD [63]. Chiral SU(3) model has been used in literature to study the modification of properties of kaons [65, 66, 64], vector mesons ($\omega,\rho$ and $\phi$) [67], pseudoscalar $\eta$ meson [68, 69], $D$ [38, 70, 39] and $B$ mesons [71], charmonium [72] and bottomonium [73], in nuclear and strange hadronic matter. The optical potentials of mesons under study, calculated using chiral effective model for different light and heavy nuclei, will be used to solve Klein Gordon equation in the momentum space to obtain the binding energies and decay width. Note that in the current work we used mean field approximation in calculating the in-medium masses and optical potentials of different mesons throughout the manuscript. In future, we may further improve the results employing beyond mean field calculations for example including two loop contribution [74, 75].
@@ II Chiral SU(3) hadronic model
To investigate the formation of mesic nuclei in the present work, we shall use the chiral SU(3) hadronic mean-field model based on nonlinear realization of chiral symmetry [63, 76, 77, 78]. In the chiral SU(3) model, the interactions between the nucleons are mediated through the exchange of non-strange scalar meson $\sigma$, strange scalar meson $\zeta$, scalar-isovector meson $\delta$, vector meson $\omega$ and vector-isovector meson $\rho$. The scalar field $\delta$ and vector field $\rho$ contribute when medium has finite isospin asymmetry. In this model, the scalar dilaton field, $\chi$, known as glueball field is introduced to incorporate the broken scale invariance property of the QCD [63]. As discussed in the introduction, QMC model, NJL model, chiral unitary approach, linear sigma model etc. have been used to calculate the in-medium optical potentials and hence the possibility of the formation of mesic nuclei. Comparing chiral SU(3) model with these approaches following are important differences. In the QMC model, quarks are treated as fundamental degrees of freedom confined inside the hadrons through a bag potential. Quarks confined inside the baryons interact through the exchange of scalar and vector fields, which in-turn modify the properties of baryons. In the QMC model in-medium masses of kaons and antikaons decrease as a function of density of the nuclear matter. Repulsive interactions for kaons and attractive for antikaons contributed by Weinberg Tomozawa term are not taken into account in this model (also for other pseudoscalar mesons). In the chiral unitary approach, coupled channel dynamics is employed to investigate the properties of antikaons and $\phi$ mesons. Coupled channel approach involve the coupling of $\bar{K}N$ to $\pi\Sigma$ and $\pi\Lambda$ channels and is important to study the dynamics of $\Lambda(1405)$ resonance also [79]. Such coupled channel effects are not considered in our present chiral model calculations. In the NJL model calculations for the $\eta$ and $\eta^{{}^{\prime}}$ mesons, positive mass-shift for $\eta$ and negative mass-shift for $\eta^{{}^{\prime}}$ were observed [8] . In the NJL model interaction Lagrangian densities are expressed in terms of quark fields.
The Lagrangian density for chiral SU(3) hadronic mean-field model is written as [63]
@@
respectively. In the above, $\gamma_{i}$ is the degeneracy factor for the nucleons. To study the properties of mesons in infinite nuclear medium at zero temperature, Eqs. (9) to (14) are solved using the expressions of $\rho_{i}$ and $\rho_{i}^{s}$ given above. For finite temperature effects, $\rho_{i}$ and $\rho_{i}^{s}$ are defined as
$\rho_{i}=\gamma_{i}\int\frac{d^{3}\mathbf{k}}{(2\pi)^{3}}\Bigg{(}\frac{1}{1+\exp\left[\beta(E^{\ast}_{i}(\mathbf{k})-\mu^{*}_{i})\right]}-\frac{1}{1+\exp\left[\beta(E^{\ast}_{i}(\mathbf{k})+\mu^{*}_{i})\right]}\Bigg{)},$ (20)
and
$\rho_{i}^{s}=\gamma_{i}\int\frac{d^{3}\mathbf{k}}{(2\pi)^{3}}\frac{m_{i}^{*}}{E^{\ast}_{i}(\mathbf{k})}\Bigg{(}\frac{1}{1+\exp\left[\beta(E^{\ast}_{i}(\mathbf{k})-\mu^{*}_{i})\right]}+\frac{1}{1+\exp\left[\beta(E^{\ast}_{i}(\mathbf{k})+\mu^{*}_{i})\right]}\Bigg{)},$ (21)
respectively. Here, $\beta=1/T$.
@@ III.1 Pseudoscalar $\eta$ and $\eta^{{}^{\prime}}$ mesons
$m_{\eta_{0}\eta_{8}}^{*}=\left[\frac{2}{\sqrt{3}\sigma_{0}c}\left(m_{\pi}^{2}f_{\pi}-m_{K}^{2}f_{K}\right)\right]^{1/2}.$ (40)
@@ III.2 Pseudoscalar $D^{0}$ and $\bar{D^{0}}$ mesons
The effective masses of neutral $D^{0}$ and $\bar{D^{0}}$ mesons required in Eq.(24) to study the $D$ mesic nuclei are obtained in the present work by extending the chiral SU(3) model to the SU(4) case [38, 84]. The mesons $D^{0}$ and $\bar{D^{0}}$ are the members of $D$ $\left(D^{+},D^{0}\right)$ and $\bar{D}$ $\left(D^{-},\bar{D^{0}}\right)$ doublets, respectively. From the Lagrangian density describing the interactions of $D$ and $\bar{D}$ mesons with nucleons in the nuclear medium, equations of motion are obtained whose Fourier transformation will give the dispersion relation
@@ III.3 Pseudoscalar $B^{0}$ and $\bar{B^{0}}$ mesons
To study the formation of bound states corresponding to $B^{0}$ and $\bar{B^{0}}$ mesons chiral SU(3) model is generalized to SU(5) sector as these bottom mesons are composed of one heavy $b$ quark/antiquark. The pseudoscalar $B^{0}$ and $\bar{B^{0}}$ mesons belong to the open bottom $B(B^{+},B^{0})$ and $\bar{B}(B^{-},\bar{B^{0}})$ meson doublets. In Ref. [43] the masses of open bottom mesons have been calculated in nuclear and strange hadronic matter using chiral effective model. For our present calculations, we need the mass shift of $B^{0}$ and $\bar{B^{0}}$ mesons in the nuclear medium at zero temperature obtained by solving the dispersion relation
@@
$+\left[-\frac{1}{f_{B}}\left(\sigma^{\prime}+\sqrt{2}\zeta_{b}^{\prime}-\delta^{\prime}\right)+\frac{d_{1}}{2f_{B}^{2}}\left(\rho_{p}^{s}+\rho_{n}^{s}\right)\right.\left.+\frac{d_{2}}{2f_{B}^{2}}\left(\rho_{p}^{s}+2\rho_{n}^{s}\right)\right]\left(\omega^{2}-\textbf{k}^{2}\right),$ (48)
@@
$+\left[-\frac{1}{f_{B}}\left(\sigma^{\prime}+\sqrt{2}\zeta_{b}^{\prime}-\delta^{\prime}\right)+\frac{d_{1}}{2f_{B}^{2}}\left(\rho_{p}^{s}+\rho_{n}^{s}\right)\right.\left.+\frac{d_{2}}{2f_{B}^{2}}\left(\rho_{p}^{s}+2\rho_{n}^{s}\right)\right]\left(\omega^{2}-\textbf{k}^{2}\right),$ (49)
@@ III.4 Pseudoscalar $\bar{K^{0}}$ mesons
$\Pi_{\bar{K^{0}}}(\omega,|\textbf{k}|,\rho(r))$ $=\frac{1}{2f_{K}^{2}}\left(\rho_{p}+2\rho_{n}\right)\omega+\frac{m_{K}^{2}}{2f_{K}}\left(\sigma^{\prime}+\sqrt{2}\zeta^{\prime}-\delta^{\prime}\right)$
$+\left[-\frac{1}{f_{K}}\left(\sigma^{\prime}+\sqrt{2}\zeta^{\prime}-\delta^{\prime}\right)+\frac{d_{1}}{2f_{K}^{2}}\left(\rho_{p}^{s}+\rho_{n}^{s}\right)\right.\left.+\frac{d_{2}}{2f_{K}^{2}}\rho_{n}^{s}\right]\left(\omega^{2}-\textbf{k}^{2}\right).$ (51)
@@ IV Results and discussion
We now discuss the possibility of the formation of mesic nuclei of various mesons through the numerical calculations of the optical potentials and binding energies. We shall calculate the optical potentials of $\eta,\eta^{{}^{\prime}}$, $D^{0}$, $\bar{D^{0}}$, $B^{0}$, $\bar{B^{0}}$,$\bar{K^{0}}$ and $\phi$ mesons in the medium relevant for the nuclei 12C, 16O, 40Ca and 208Pb. As we discussed in the previous section, the in-medium properties of mesons are calculated using the chiral SU(3) hadronic mean field model and extended to SU(4) and SU(5) sector to obtain the in-medium masses of charmed and bottom mesons [63]. The parameters $k_{0}$, $k_{2}$ and $k_{4}$ of chiral SU(3) model appearing in Eq. 5 are fitted to reproduce the vacuum values of $\sigma,\zeta$ and $\chi$ fields [63]. Further, the vacuum values $\sigma_{0}$ and $\zeta_{0}$ are calculated using the values of pion decay constant $f_{\pi}=93.3$ and kaon decay constant $f_{K}=122$ MeV through relations $\sigma_{0}=-f_{\pi}$ and $\zeta_{0}=-\frac{1}{\sqrt{2}}(2f_{K}-f_{\pi})$. The vacuum value of $\chi_{0}$ is constrained so that the binding energy $-16$ MeV for symmetric nuclear matter at saturation density $\rho_{0}=0.15$ fm-3 is reproduced. Also, the parameter $k_{3}$ is constrained through the mass of $\eta$ and $\eta^{{}^{\prime}}$ meson and $k_{1}$ is fixed to give a mass of the order of $m_{\sigma}\sim$ $500$ MeV. The values of parameters, $d_{1}=2.56/m_{K}$ and $d_{2}=0.73/m_{K}$ are fitted to empirical values of kaon-nucleon scattering length [90, 65, 66].
@@
Figure 1: The mass shift of $\eta$ [in subplots (a) and (c)] and $\eta^{\prime}$ [in subplots (b) and (d)] mesons are shown as a function of density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of nuclear medium for isospin asymmetry, $I=0,0.211$ and $0.3$. Results are shown for temperatures $T=0$ and $100$ MeV.
@@ IV.1 Binding energies for $\eta$ and $\eta^{{}^{\prime}}$ mesic nuclei
In this subsection, we shall explore the possibility of the formation of $\eta$ and $\eta^{{}^{\prime}}$ mesic nuclei. Using Eq. (41), the mass shifts of pseudoscalar $\eta$ and $\eta^{{}^{\prime}}$ mesons are calculated within the chiral SU(3) model. In Fig. 1, we have shown the mass shift of $\eta$ and $\eta^{{}^{\prime}}$ mesons as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of the infinite nuclear medium for isospin asymmetry $I=0,0.211$ and $0.3$, at temperatures $T=0$ and $100$ MeV. The isospin asymmetry $I=0.211$ corresponds to the heavy nuclei 208Pb considered in the present work for the bound state formation. As we discussed earlier, for the calculations of $\eta$ and $\eta^{{}^{\prime}}$ meson properties, we considered the mixing of $\eta_{8}$ and $\eta_{0}$ states. The pseudoscalar $\eta$ meson which is member of octet is observed to undergo larger mass shift as compared to the singlet $\eta^{{}^{\prime}}$. At nuclear saturation density $\rho_{B}=\rho_{0}$, for isospin asymmetry $I=0(0.3)$ and temperature $T=0$ MeV, the values of mass shift $\Delta m_{\eta}$ and $\Delta m_{\eta}^{{}^{\prime}}$ are observed to be $-34.64(-34.21)$ and $-1.03(-1.02)$ MeV, respectively. At baryon density $\rho_{B}=4\rho_{0}$, the above values of mass shift change to $-124.15(-120.92)$ and $-3.00(-2.94)$ MeV. The finite isopsin asymmetry of the medium causes less drop in the masses of $\eta$ and $\eta^{{}^{\prime}}$, although the impact is small. An increase in the temperature of nuclear medium from $T=0$ to $100$ MeV, enhances the in-medium mass of these mesons. Furthermore, in the present chiral model calculations, compared to $\eta$ mesons, the singlet $\eta^{{}^{\prime}}$ undergoes very less drop in mass as a function of density of nuclear medium. Recall that the in-medium masses of $\eta$ and $\eta^{{}^{\prime}}$ mesons are obtained from the diagonalization of Eq. (41). Neglecting the off-diagonal terms (having very small contributions), $m_{\eta}^{*}\sim m_{\eta_{8}}^{*}$ and $m_{\eta{{}^{\prime}}}^{*}\sim m_{\eta_{0}}^{*}$. As can be seen from Eq.(39), within the chiral SU(3) model $m_{\eta_{0}}^{*}=m_{\eta_{0}}$, i.e., $\eta_{0}$ is not modified in the medium and therefore, $\eta^{{}^{\prime}}$ mesons are not modified much in the present calculations (small modifications are stimulated through the diagonalization of Eq. (41).
@@ IV.2 Binding energies for $D^{0}$ and $\bar{D^{0}}$ mesic nuclei
Now we discuss the formation of bound states for neutral $D^{0}$ and $\bar{D^{0}}$ mesons with the nuclei within the chiral SU(3) hadronic mean field model. The mass shift of $D^{0}$ and $\bar{D^{0}}$ mesons calculated using Eq. (46) are plotted in Fig. (3) as a function of baryon density ratio $\rho_{B}/\rho_{0}$. As can be seen from the figure, the in-medium masses of both $D^{0}$ and $\bar{D^{0}}$ mesons decrease as a function of density of the nuclear medium. Furthermore, the increase of isospin asymmetry of the medium from zero to finite value (which is relevant for the study of nuclei having different number of protons and neutrons) decreases the magnitude of mass shift for these neutral pseudoscalar mesons. At zero temperature and baryon density $\rho_{B}=\rho_{0}(4\rho_{0})$, the values of mass shift for $D^{0}$ and $\bar{D^{0}}$ mesons are observed to $-76.96(-343.82)$ and $-26(-159.71)$ MeV in the symmetric nuclear medium, whereas for $I=0.3$ these values changes to $-64.17(-298.80)$ and $-23.37(-148.43)$ MeV, respectively. For temperature $T=100$ MeV, the values of mass shift for $D^{0}$ and $\bar{D^{0}}$ mesons are observed to be $-64.92(-316.43)$ and $-14.24(-129.53)$ MeV, respectively, in symmetric nuclear medium at $\rho_{B}=\rho_{0}(4\rho_{0})$.
Using the observed values of mass shift for $D^{0}$ and $\bar{D^{0}}$ mesons at center of given nuclei in Eq. (24), the real part of optical potential is calculated and is plotted in Fig. 4 as a function of $r$ for nuclei, 12C, 16O, 40Ca and 208Pb. As one can see from the figure, depth of optical potential (large negative value) is observed to be more for $D^{0}$ mesons as compared to $\bar{D^{0}}$. In case of $\bar{D^{0}}$ mesons the optical potential $U_{\bar{D^{0}}}(r)$ is observed to have positive values for certain range of $r$ causing a bump as can be seen from Fig. 4(b). On the otherside, such positive values are not observed for the optical potential of $D^{0}$ mesons. The observed difference is because of opposite contribution of Weinberg Tomozawa term for the $D^{0}$ and $\bar{D^{0}}$ mesons as can be seen from Eqs. (44) and (45), respectively. Weinberg Tomozawa term alone gives positive contribution to the optical potential of $\bar{D^{0}}$ mesons, whereas the combined contribution of all other terms give negative value for $U_{\bar{D^{0}}}(r)$. For certain range of $r$, the positive contribution due to Weinberg Tomozawa term dominates over the net negative contribution of other terms, causing overall positive value of optical potential for $\bar{D^{0}}$ mesons. To understand this in Fig. 4(c) and Fig. 4(d), we plot the optical potentials of $D^{0}$ and $\bar{D^{0}}$ mesons, respectively, considering (i) Weinberg Tomozawa term only and (ii) without Weinberg Tomozawa term, for 208Pb. As can be seen from Fig. 4(c), in case of $D^{0}$ mesons optical potentials from Weinberg term have negative values and therefore, unlike $\bar{D^{0}}$ mesons, positive values are not observed for $D^{0}$ mesons.
The imaginary part of optical potential, required to understand the absorption of mesons in the nuclei and its impact on the binding energy of neutral $D$ mesic nuclei, is calculated using Eq. (25). The values of binding energy and decay width obtained by solving the Klein Gordon equation are given in Tables 2 and 3, for $D^{0}$ and $\bar{D^{0}}$ mesons, respectively, for $\kappa=0,0.5$ and 1. As one can see from Table 2, the bound states of $D^{0}$ mesons, for 40Ca and 208Pb may be found not only for lower states $1s$ and $1p$, but also for other higher states when $\kappa=0$ and $0.5$. When the value of $\kappa$ is increased to one, the magnitude of binding energy is observed to be less as compared to the decay width value. This may cause a broadening of the peak and hindrance in the observation of clear signature of bound states. Comparing the values of binding energies in Tables 2 and 3, magnitude of binding energies for $\bar{D^{0}}$ mesons are significantly less than $D^{0}$ mesons. In Ref. [10], the bound states of $D^{0}$ mesons have been studied for several nuclei using the coupled channel approach. Comparing our results for 12C, 40Ca and 208Pb with Ref. [10], the magnitude of binding energies are observed to be more in our case. The binding energies of $D^{0}$ and $\bar{D^{0}}$ mesons for 208Pb nuclei are calculated using QMC model and are found to approximately $-96$ and $-25$ MeV, respectively [7].
@@
Figure 3: The mass shift of $D^{0}$ [in subplots (a) and (c)] and $\bar{D^{0}}$ [in subplots (b) and (d)] mesons are shown as a function of baryon density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of nuclear medium for isospin asymmetry, $I=0,0.211$ and $0.3$. Results are shown for temperatures $T=0$ and $100$ MeV.
@@
Figure 4: The real optical potentials $U(r)$ of $D^{0}$ [in subplot (a)] and $\bar{D^{0}}$ [in subplot (b)] mesons are shown as a function of distance $r$ from center, for nuclei 12C, 16O, 40Ca and 208Pb, with mass number, $A=12,16,40$ and $208$, respectively. In subplots (c) and (d) optical potentials are plotted for $A=208$ considering the impact of Weinberg Tomozawa (WT) term to the total optical potential.
@@
Figure 5: The mass shift of $B^{0}$ [in subplots (a) and (c)] and $\bar{B^{0}}$ [in subplots (b) and (d)] mesons are shown as a function of density $\rho_{B}$ (in units of nuclear saturation density $\rho_{0}$) of nuclear medium for isospin asymmetry, $I=0,0.211$ and $0.3$. Results are shown for temperatures $T=0$ and $100$ MeV.